# Initial value space of the fourth order Painlevé system with $A_{5}^{(1)}\times A_{1}^{(1)}$ symmetry

### v2

@@ 5 Resolving the accessible singularities $P_{j}^{(kl)}$
Coordinates resolving the accessible singularities (3.4) are given as follows.
@@
In this proposition, we denote the coordinates resolving $P_{j}^{(kl)}$ by $\left(\widetilde{x}_{j}^{[kl]},\widetilde{y}_{j}^{[kl]},\widetilde{z}_{j}^{[kl]},\widetilde{w}_{j}^{[kl]}\right)$. We also use the notation
$r_{3}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ $=\left(r_{3}\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1}\right),r_{3}(q_{2}p_{1}),r_{3}\left(\dfrac{1}{p_{1}}\right),r_{3}\left(\dfrac{p_{2}}{p_{1}}\right)\right)$
$=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},\frac{(q_{2}p_{2}+\alpha_{3})p_{1}}{p_{2}},\frac{1}{p_{1}},\frac{p_{2}}{p_{1}}\right)$
for the sake of simplicity. Recall that the coordinate $\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$ was given in Proposition 4.1 as
$\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right).$
@@
Hence the solution to system (4.1) with initial condition (5.1) is transformed to that to the system in terms of the coordinate $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1}^{(01)},\widetilde{p}_{2}^{(01)}\right)$ whose initial condition is given by
@@
Since system (1.1) is invariant under the transformation $r_{3}$, the system in terms of $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1},\widetilde{p}_{2}\right)$ is the same as system (1.1) except changing the parameters. Hence the system in terms of $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1}^{(01)},\widetilde{p}_{2}^{(01)}\right)$ is the same as system (4.1) except changing the parameters and has the accessible singularity
@@
It is obvious that point (5.8) is contained in accessible singularity (5.9). On the other hand, we have given the coordinate resolving
@@
and it is a symplectic transformation. Note that the system in terms of $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ is described as a polynomial Hamiltonian system.
In the last, we see the solution to the system in terms of $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$. Relations (5.7) and (5.10) imply
@@
where $A_{1},\ldots,A_{5}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta]$. Hence the solution to system (4.1) with initial condition (5.1) is transformed to that in terms of $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ whose initial condition is given by
@@ 6 Main result
We have given 9 and 12 coordinates resolving accessible singularities (3.3) and (3.4) in Proposition 4.1 and 5.1 respectively. Hence we can construct an initial value space of system (1.1).
@@
An initial value space of system (1.1) is constructed by glueing 24 coordinates with relations
@@
$\left(\widetilde{x}_{7}^{[22]},\widetilde{y}_{7}^{[22]},\widetilde{z}_{7}^{[22]},\widetilde{w}_{7}^{[22]}\right)$ $=r_{3}r_{2}\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right),$
$\left(\widetilde{x}_{4}^{[22]},\widetilde{y}_{4}^{[22]},\widetilde{z}_{4}^{[22]},\widetilde{w}_{4}^{[22]}\right)$ $=r_{2}r_{1}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right).$
@@
We also consider the holomorphy property of system (1.1). Choose 6 from 9 coordinates given in Proposition 4.1 and set
$\begin{split}(x_{0},y_{0},z_{0},w_{0})&=\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)\\ &=\left(-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{2}\right),\\ (x_{1},y_{1},z_{1},w_{1})&=\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right)\\ &=\left(-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{1}^{(1)},q_{2}^{(1)}p_{1}^{(1)},\dfrac{1}{p_{1}^{(1)}},\dfrac{p_{2}^{(1)}}{p_{1}^{(1)}}\right),\\ (x_{2},y_{2},z_{2},w_{2})&=\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right)\\ &=\left(q_{1}^{(1)},p_{1}^{(1)},-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},\dfrac{1}{p_{2}^{(1)}}\right),\\ (x_{3},y_{3},z_{3},w_{3})&=\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right)\\ &=\left(-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{1}^{(2)},q_{2}^{(2)}p_{1}^{(2)},\dfrac{1}{p_{1}^{(2)}},\dfrac{p_{2}^{(2)}}{p_{1}^{(2)}}\right),\\ (x_{4},y_{4},z_{4},w_{4})&=\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right)\\ &=\left(q_{1}^{(2)},-\left\{\left(q_{2}^{(2)}-\frac{1}{t}\right)p_{2}^{(2)}-\alpha_{4}\right\},p_{1}^{(2)},\frac{1}{p_{2}^{(2)}}p_{2}^{(2)}\right),\\ (x_{5},y_{5},z_{5},w_{5})&=\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)\\ &=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right).\end{split}$ (6.1)
Then we obtain the following proposition.
Proposition 6.3.
There exists a unique polynomial Hamiltonian system of degree 5 with respect to $q_{1},q_{2},p_{1},p_{2}$ which is holomorphic on each coordinates $(x_{i},y_{i},z_{i},w_{i})\ (i=0,\ldots,5)$. Moreover, this Hamiltonian system coincides with (1.1).
Proof.
Let $K$ be a Hamiltonian of degree 5 described as
$K=\sum_{k=1}^{5}\sum_{\begin{subarray}{c}j_{1}+j_{2}+j_{3}+j_{4}=k,\\[1.0pt] j_{1}\geq 0,\,j_{2}\geq 0,\,j_{3}\geq 0,\,j_{4}\geq 0\end{subarray}}a_{j_{1},j_{2},j_{3},j_{4}}q_{1}^{j_{1}}q_{2}^{j_{2}}p_{1}^{j_{3}}p_{2}^{j_{4}}.$ (6.2)
We first consider the coordinate $(x_{0},y_{0},z_{0},w_{0})$. The first relation of (6.1) is rewritten to
$q_{1}=1+(\alpha_{0}-x_{0}z_{0})z_{0},\quad q_{2}=y_{0},\quad p_{1}=\frac{1}{z_{0}},\quad p_{2}=w_{0}.$ (6.3)
Substituting relation (6.3) to (6.2), we obtain
$K=\frac{a_{0,0,5,0}}{z_{0}^{5}}+\frac{F_{1}}{z_{0}^{4}},$
where $F_{1}\in\mathbb{C}[x_{0},y_{0},z_{0},w_{0}]$. Then the coefficient $a_{0,0,5,0}$ is determined uniquely as $a_{0,0,5,0}=0$ and the Hamiltonian $K$ is descried as
$K=\frac{a_{0,0,4,0}+a_{1,0,4,0}+a_{0,1,4,0}y_{0}+a_{0,0,4,1}w_{0}}{z_{0}^{4}}+\frac{F_{2}}{z_{0}^{3}},$
where $F_{2}\in\mathbb{C}[x_{0},y_{0},z_{0},w_{0}]$. It follows that
$a_{0,0,4,0}+a_{1,0,4,0}=0,\quad a_{0,1,4,0}=0,\quad a_{0,0,4,1}=0.$
We repeat this procedure until $K$ is a polynomial with respect to $z_{0}$ and do the same for the other coordinates $(x_{i},y_{i},z_{i},w_{i})\ (i=1,\ldots,5)$. Then all of the coefficients $a_{j_{1},j_{2},j_{3},j_{4}}$ are determined uniquely and the Hamiltonian $K$ coincides with that of system (1.1). Note that we have to consider the Hamiltonian $K+\frac{p_{2}^{(2)}}{t^{2}}$ for the coordinate $(x_{4},y_{4},z_{4},w_{4})$ because the independent variable $t$ appears in the fifth relation of (6.1). ∎
@@
Sasano considered the holomorphy property of system (1.1) in terms of 7 coordinates in [13]. The number of coordinates in Proposition 6.3 is 1 less than that in [13] because we always assume that $\alpha_{0}+\cdots+\alpha_{5}=1$.

### v3

@@ 6 Resolving the accessible singularities $P^{0}$
Coordinates resolving the accessible singularity $P^{0}$ are given as follows.
@@
Hence the solution to system (3.6) with initial condition (6.1) is transformed to that to the system on the coordinate $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1}^{(01)},\widetilde{p}_{2}^{(01)}\right)$ whose initial condition is given by
@@
Since system (2.2) is invariant under the transformation $r_{3}$, the system on $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1},\widetilde{p}_{2}\right)$ is the same as system (2.2) except changing the parameters. Hence the system on $\left(\widetilde{q}_{1},\widetilde{q}_{2},\widetilde{p}_{1}^{(01)},\widetilde{p}_{2}^{(01)}\right)$ is the same as system (3.6) except changing the parameters and has the accessible singularity
@@
It is obvious that point (6.8) is contained in accessible singularity (6.9). Note that we have
$\left\{\widetilde{q}_{1}=\widetilde{q}_{2}=\widetilde{p}_{1}^{(01)}=0,\ \widetilde{p}_{2}^{(01)}\neq 0\right\}=\left\{q_{1}=q_{2}=p_{1}^{(01)}=0,\ p_{2}^{(01)}\neq 0\right\}.$
On the other hand, we have given the coordinate resolving
@@
and it is a symplectic transformation. Note that the system on $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ is described as a polynomial Hamiltonian system.
In the last, we see the solution to the system on $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$. Relations (6.7) and (6.10) imply
@@
where $A_{1},\ldots,A_{5}\in\mathbb{C}(t_{0})[\alpha_{0},\ldots,\alpha_{5},\eta]$. Hence the solution to system (3.6) with initial condition (6.1) is transformed to that on $\left(\widetilde{x}_{3}^{[01]},\widetilde{y}_{3}^{[01]},\widetilde{z}_{3}^{[01]},\widetilde{w}_{3}^{[01]}\right)$ whose initial condition is given by
@@ 7 Main result
We summarize the facts obtained in the previous sections. The system $\mathcal{H}$ has the singularity $D^{0}$ given in Proposition 3.5. There exists the accessible singularity $C^{0}\cup P^{0}$ on $D^{0}$ given in (4.2). All of leaves corresponding the solutions passing through the singularity $D^{0}\setminus\left(C^{0}\cup P^{0}\right)$ are vertical leaves.
Consider the first blowing up along the accessible singularity $C^{0}\cup P^{0}$ and denote the proper image of $D^{0}$ by the same symbol. Then we obtain the exceptional divisor
$D^{1}$ $=D_{1}^{1}\cup D_{2}^{1}\cup D_{3}^{1}\cup r_{0}\left(D_{1}^{1}\right)\cup r_{1}r_{0}\left(D_{1}^{1}\right)\cup r_{3}\left(D_{1}^{1}\right)\cup r_{4}r_{3}\left(D_{1}^{1}\right)$
$\quad\cup r_{5}\left(D_{2}^{1}\right)\cup r_{0}r_{5}\left(D_{2}^{1}\right)\cup r_{2}\left(D_{2}^{1}\right)\cup r_{3}r_{2}\left(D_{2}^{1}\right)\cup r_{4}\left(D_{3}^{1}\right)\cup r_{5}r_{4}\left(D_{3}^{1}\right)\cup r_{1}\left(D_{3}^{1}\right)\cup r_{2}r_{1}\left(D_{3}^{1}\right).$
Here the component $D_{1}^{1}$ is given in (5.2). The components $D_{2}^{1}$ and $D_{3}^{1}$ are obtained by the blowing up along $C_{2}^{0}$ and $C_{3}^{0}$ respectively in a similar manner as $D_{1}^{1}$. Recall that the accessible singularities $C_{2}^{0}$ and $C_{3}^{0}$ are given in Lemma 4.4. The rest 12 components are obtained by the actions of the transformations $r_{0},\ldots,r_{5}$ as is seen in Section 6. The symbol $r_{0}\left(D_{1}^{1}\right)$ stands for a set given as follows. If we choose a component $\left\{v_{11}=0\right\}$ of $D_{1}^{1}$, then we have
$r_{0}\left(\{v_{11}=0\}\right)=\left\{r_{0}(v_{11})=0\right\}=\left\{r_{0}\left(p_{1}^{(01)}\right)=0\right\}.$
There exists the accessible singularity $C^{1}\cup P^{1}$ with
$C^{1}$ $=C_{1}^{1}\cup C_{2}^{1}\cup C_{3}^{1}\cup r_{0}\left(C_{1}^{1}\right)\cup r_{1}r_{0}\left(C_{1}^{1}\right)\cup r_{3}\left(C_{1}^{1}\right)\cup r_{4}r_{3}\left(C_{1}^{1}\right)$
$\quad\cup r_{5}\left(C_{2}^{1}\right)\cup r_{0}r_{5}\left(C_{2}^{1}\right)\cup r_{2}\left(C_{2}^{1}\right)\cup r_{3}r_{2}\left(C_{2}^{1}\right)\cup r_{4}\left(C_{3}^{1}\right)\cup r_{5}r_{4}\left(C_{3}^{1}\right)\cup r_{1}\left(C_{3}^{1}\right)\cup r_{2}r_{1}\left(C_{3}^{1}\right)$
and
$P^{1}=P_{1}^{1}\cup P_{2}^{1}\cup P_{3}^{1}.$
Here the component $C_{1}^{1}\cup P_{1}^{1}$ is given in (5.8). The components $C_{2}^{1}\cup P_{2}^{1}$ and $C_{3}^{1}\cup P_{3}^{1}$ are the accessible singularities on $D_{2}^{1}$ and $D_{3}^{1}$ respectively. All of leaves corresponding the solutions passing through the singularity $D^{1}\setminus\left(C^{1}\cup P^{1}\right)$ are vertical leaves.
Consider the second blowing up along the accessible singularity $C^{1}\cup P^{1}$ and denote the proper image of $D^{1}$ by the same symbol. Then we obtain the exceptional divisor
$D^{2}$ $=D_{1}^{2}\cup r_{2}\left(D_{1}^{2}\right)\cup D_{2}^{2}\cup r_{4}\left(D_{2}^{2}\right)\cup D_{3}^{2}\cup r_{0}\left(D_{3}^{2}\right)\cup r_{0}\left(D_{1}^{2}\right)\cup r_{1}r_{0}\left(D_{1}^{2}\right)\cup r_{3}\left(D_{1}^{2}\right)\cup r_{4}r_{3}\left(D_{1}^{2}\right)$
$\quad\cup r_{5}\left(D_{2}^{2}\right)\cup r_{0}r_{5}\left(D_{2}^{2}\right)\cup r_{2}\left(D_{2}^{2}\right)\cup r_{3}r_{2}\left(D_{2}^{2}\right)\cup r_{4}\left(D_{3}^{2}\right)\cup r_{5}r_{4}\left(D_{3}^{2}\right)\cup r_{1}\left(D_{3}^{2}\right)\cup r_{2}r_{1}\left(D_{3}^{2}\right).$
Here the component $D_{1}^{2}$ is given in (5.9). The components $D_{2}^{2}$ and $D_{3}^{2}$ are obtained by the blowing up along $C_{2}^{1}$ and $C_{3}^{1}$ respectively in a similar manner as $D_{1}^{2}$. The rest 15 components are obtained by the actions of the transformations $r_{0},\ldots,r_{5}$. There is no singularity on $D^{2}\setminus D^{1}$ and all of leaves corresponding the solutions passing through the singularity $D^{2}\cap D^{1}$ are vertical leaves.
$D^{0}$$D_{1}^{1}$$D_{1}^{2}$$r_{2}\left(D_{1}^{2}\right)$$D_{2}^{1}$$D_{2}^{2}$$r_{4}\left(D_{2}^{2}\right)$$D_{3}^{1}$$D_{3}^{2}$$r_{0}\left(D_{3}^{2}\right)$$r_{0}\big(D_{1}^{1}\big)$$r_{0}\big(D_{1}^{2}\big)$$r_{2}r_{1}\big(D_{3}^{1}\big)$$r_{2}r_{1}\big(D_{3}^{2}\big)\cdots\cdots$
Figure 1: Exceptional divisors
We illustrate the above blowing up procedure in Figure 1. Consequently, a compact manifold $\overline{E}$ is derived from that $\Sigma_{\eta}\times B$. Then we obtain an initial value space $E=\overline{E}\setminus\left(D^{0}\cup D^{1}\right)$ of system (2.2).
@@
The manifold $E$ is constructed by glueing 33 coordinates with relations
@@
$\left(x_{1}^{[01]},y_{1}^{[01]},z_{1}^{[01]},w_{1}^{[01]}\right)$ $=\left(-\{(q_{1}-q_{2})p_{1}-\alpha_{2}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{1}+p_{2}\right),$
$\left(x_{1}^{[11]},y_{1}^{[11]},z_{1}^{[11]},w_{1}^{[11]}\right)$ $=\left(-\left\{\left(q_{1}^{(1)}-\dfrac{q_{2}^{(1)}}{t}\right)p_{1}^{(1)}-\alpha_{4}\right\}p_{1}^{(1)},q_{2}^{(1)},\dfrac{1}{p_{1}^{(1)}},p_{2}^{(1)}+\dfrac{p_{1}^{(1)}}{t}\right),$
@@
$\left(x_{1}^{[21]},y_{1}^{[21]},z_{1}^{[21]},w_{1}^{[21]}\right)$ $=\left(-\left\{\left(q_{1}^{(2)}-q_{2}^{(2)}\right)p_{1}^{(2)}-\alpha_{0}\right\}p_{1}^{(2)},q_{2}^{(2)},\dfrac{1}{p_{1}^{(2)}},p_{1}^{(2)}+p_{2}^{(2)}\right),$
$\left(\widehat{x}_{3}^{[01]},\widehat{y}_{3}^{[01]},\widehat{z}_{3}^{[01]},\widehat{w}_{3}^{[01]}\right)$ $=r_{2}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right),$
$\left(\widehat{x}_{3}^{[02]},\widehat{y}_{3}^{[02]},\widehat{z}_{3}^{[02]},\widehat{w}_{3}^{[02]}\right)$ $=r_{2}\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right),$
$\left(\widehat{x}_{3}^{[11]},\widehat{y}_{3}^{[11]},\widehat{z}_{3}^{[11]},\widehat{w}_{3}^{[11]}\right)$ $=r_{4}\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right),$
$\left(\widehat{x}_{3}^{[12]},\widehat{y}_{3}^{[12]},\widehat{z}_{3}^{[12]},\widehat{w}_{3}^{[12]}\right)$ $=r_{4}\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right),$
$\left(\widehat{x}_{3}^{[21]},\widehat{y}_{3}^{[21]},\widehat{z}_{3}^{[21]},\widehat{w}_{3}^{[21]}\right)$ $=r_{2}\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right),$
$\left(\widehat{x}_{3}^{[22]},\widehat{y}_{3}^{[22]},\widehat{z}_{3}^{[22]},\widehat{w}_{3}^{[22]}\right)$ $=r_{2}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right),$
@@
$\left(\widetilde{x}_{4}^{[22]},\widetilde{y}_{4}^{[22]},\widetilde{z}_{4}^{[22]},\widetilde{w}_{4}^{[22]}\right)$ $=r_{2}r_{1}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right),$
@@
$\left(\widetilde{x}_{7}^{[22]},\widetilde{y}_{7}^{[22]},\widetilde{z}_{7}^{[22]},\widetilde{w}_{7}^{[22]}\right)$ $=r_{3}r_{2}\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right).$
This theorem follows from Propositions 5.1 and 6.1.
@@
There exists a unique polynomial Hamiltonian system of degree 5 with respect to $q_{1},q_{2},p_{1},p_{2}$ which is holomorphic on each coordinates $(x_{i},y_{i},z_{i},w_{i})\ (i=0,\ldots,5)$. Moreover, this Hamiltonian system coincides with (2.2). This fact was first mentioned by Sasano in [17]. It is expected that we can show uniqueness in a similar manner as in [19]. However, we have not succeeded in the proof yet.